Consider this aqueous reaction. HNO,(a)+ Ba(OH) aq)> What is the formula for the salt that forms?
Answer #1

The balanced reaction

HNO3(aq) + Ba(OH)2(aq) = Ba(NO32)2(s) + 2H2O(l)

Nitric acid + Barium hydroxide = Barium nitrate + water

Formula of Salt forms = Ba(NO32)2